# CuClarabel: GPU Acceleration for a Conic Optimization Solver

### v2

@@ Appendix D Detailed benchmark results
case2383wp_k 97600 20393 155950 0 109 - 107 - 1.81 - 3.63 -
case2383wp_k__api 97600 20393 155950 0 21 38 23 33 0.928 1.21 1.26 1.89
case2383wp_k__sad 97600 20393 155950 0 99 - 99 - 1.78 - 3.48 -
case2312_goc 98556 20518 159516 0 47 151 48 - 1.35 3.98 2.05 -
case2312_goc__api 98556 20518 159516 0 65 250 65 - 1.41 6.23 2.19 -
case2312_goc__sad 98556 20518 159516 0 59 191 59 - 1.32 4.98 1.98 -
case2737sop_k 109822 22777 176602 0 70 - 75 - 1.6 - 3.07 -
case2737sop_k__api 109822 22777 176602 0 71 - 73 - 1.63 - 2.99 -
case2737sop_k__sad 109822 22777 176602 0 72 - 73 - 1.62 - 2.92 -
case2736sp_k 110022 22878 176901 0 71 - 69 - 1.63 - 2.88 -
case2736sp_k__api 110022 22878 176901 0 57 83 57 - 1.41 2.63 2.65 -
case2736sp_k__sad 110022 22878 176901 0 71 - 72 - 1.53 - 3.67 -
case2746wp_k 111106 23320 178556 0 74 - 73 - 1.78 - 3.28 -
case2746wp_k__api 111106 23320 178556 0 31 33 32 35 1.18 1.22 1.78 2.36
case2746wp_k__sad 111106 23320 178556 0 74 133 74 - 1.65 4.28 3.19 -
case2746wop_k 111822 23434 179829 0 64 - 64 - 1.57 - 2.8 -
case2746wop_k__api 111822 23434 179829 0 26 33 - 36 1.11 1.18 - 2.21
case2746wop_k__sad 111822 23434 179829 0 64 - 63 - 1.43 - 2.6 -
case3012wp_k 120676 25202 193907 0 94 - 91 - 1.96 - 3.94 -
case3012wp_k__api 120676 25202 193907 0 29 53 - 48 1.2 1.98 - 3.1
case3012wp_k__sad 120676 25202 193907 0 89 - 91 - 1.94 - 3.92 -
case2848_rte 122708 25858 197228 0 80 177 80 - 2.15 5.3 4.75 -
case2848_rte__api 122708 25858 197228 0 75 - 75 - 2.28 - 5.25 -
case2848_rte__sad 122708 25858 197228 0 70 - 70 - 2.39 - 4.36 -
case2868_rte 123912 26164 198869 0 84 225 84 - 2.35 6.49 5.8 -
case2868_rte__api 123912 26164 198869 0 73 258 75 - 2.11 7.53 4.96 -
case2868_rte__sad 123912 26164 198869 0 79 288 79 - 2.24 8.31 5.24 -
case3120sp_k 124354 25856 199827 0 79 - 75 - 1.77 - 3.4 -
case3120sp_k__api 124354 25856 199827 0 87 - 92 - 2.02 - 4.14 -
case3120sp_k__sad 124354 25856 199827 0 99 - 100 - 2.03 - 4.4 -
case2853_sdet 128886 27445 206896 0 102 269 103 - 2.1 8.51 4.34 -
case2853_sdet__api 128886 27445 206896 0 122 309 121 - 2.43 9.57 5.12 -
case2853_sdet__sad 128886 27445 206896 0 101 276 101 - 2.1 8.64 4.33 -
case3022_goc 134780 28060 217434 0 41 174 - 139 1.52 6.16 - 9.67
case3022_goc__api 134780 28060 217434 0 58 202 - 177 1.84 6.88 - 12.9
case3022_goc__sad 134780 28060 217434 0 42 177 - 142 1.51 5.82 - 9.91
case3375wp_k 139126 29112 223999 0 101 - 89 - 2.14 - 4.04 -
case3375wp_k__api 139126 29112 223999 0 146 - 151 - 2.81 - 6.84 -
case3375wp_k__sad 139126 29112 223999 0 103 - 92 - 2.17 - 4.25 -
case2869_pegase 143608 30153 236217 0 53 235 53 - 2.03 8.56 4.44 -
case2869_pegase__api 143608 30153 236217 0 50 264 50 - 1.88 9.44 4.4 -
case2869_pegase__sad 143608 30153 236217 0 48 278 48 - 1.86 9.99 4.08 -
case2742_goc 144110 29856 236467 0 110 119 109 - 3.89 4.9 9.42 -
case2742_goc__api 144110 29856 236467 0 104 127 102 129 3.46 5.25 8.19 10.4
case2742_goc__sad 144110 29856 236467 0 107 119 106 - 3.83 4.94 9.09 -
case4661_sdet 198498 41599 320728 0 140 - 137 - 4.23 - 11.2 -
case4661_sdet__api 198498 41599 320728 0 125 - 125 - 3.94 - 10.2 -
case4661_sdet__sad 198498 41599 320728 0 139 - 134 - 4.37 - 10.9 -
case3970_goc 205819 42789 337328 0 140 154 143 - 4.94 9.11 15.6 -
case3970_goc__api 205819 42789 337328 0 109 136 115 - 4.85 8.04 15.1 -
case3970_goc__sad 205819 42789 337328 0 142 169 144 - 5.07 10.2 16.2 -
case4020_goc 216455 44877 355706 0 181 194 186 - 6.86 12.1 25.1 -
case4020_goc__api 216455 44877 355706 0 101 144 101 - 4.77 9.28 15.4 -
case4020_goc__sad 216455 44877 355706 0 183 178 185 - 6.71 13.2 24.4 -
case4917_goc 218213 45522 352833 0 52 164 - 147 2.99 8.63 - 16.7
case4917_goc__api 218213 45522 352833 0 62 260 60 - 3.27 13.4 7.7 -
case4917_goc__sad 218213 45522 352833 0 47 184 - 166 2.91 10.1 - 18.8
case4601_goc 225588 46885 368445 0 179 192 180 - 5.98 11.5 21.8 -
case4601_goc__api 225588 46885 368445 0 103 147 109 162 4.66 8.99 15.6 20.5
case4601_goc__sad 225588 46885 368445 0 180 191 177 - 5.94 12.2 20 -
case4837_goc 240160 49855 392884 0 139 152 137 - 7.18 10.6 24.3 -
case4837_goc__api 240160 49855 392884 0 122 177 124 - 5.95 12.1 19 -
case4837_goc__sad 240160 49855 392884 0 136 161 130 - 7.05 11 23 -
case4619_goc 254753 52616 419210 0 145 153 144 - 7.98 11 26.8 -
case4619_goc__api 254753 52616 419210 0 130 161 130 - 6.85 11.9 21.4 -
case4619_goc__sad 254753 52616 419210 0 151 150 151 - 8.17 11.2 27 -
case5658_epigrids 282180 58620 461724 0 150 181 143 - 9.01 13.8 28.4 -
case5658_epigrids__api 282180 58620 461724 0 164 202 159 - 10.8 15 34.9 -
case5658_epigrids__sad 282180 58620 461724 0 156 221 158 - 8.82 16.6 29.1 -
case6468_rte 286248 59396 463599 0 104 250 101 - 5.5 16.1 17.7 -
case6468_rte__api 286248 59396 463599 0 88 292 88 - 5.43 18.4 17.2 -
case6468_rte__sad 286248 59396 463599 0 93 254 99 - 5.2 16.5 17.5 -
case6470_rte 287806 60144 465757 0 122 319 109 - 6.14 20.3 21.9 -
case6470_rte__api 287806 60144 465757 0 86 330 85 - 5.22 21.4 17.2 -
case6470_rte__sad 287806 60144 465757 0 102 350 99 - 5.49 22.2 17.5 -
case6495_rte 288050 60099 465939 0 112 329 115 - 6.07 21.3 20.1 -
case6495_rte__api 288050 60099 465939 0 102 349 95 - 5.61 22.3 16.6 -
case6495_rte__sad 288050 60099 465939 0 112 344 108 - 5.64 22.1 18.7 -
case6515_rte 288710 60239 466922 0 97 299 95 - 6.3 19.1 17.7 -
case6515_rte__api 288710 60239 466922 0 81 332 81 - 5.29 21.7 15.8 -
case6515_rte__sad 288710 60239 466922 0 86 327 84 - 5.24 21.4 15.6 -
case7336_epigrids 358450 74618 585766 0 156 196 154 - 11.6 18 47.8 -
case7336_epigrids__api 358450 74618 585766 0 150 214 152 - 10.5 19.9 43.6 -
case7336_epigrids__sad 358450 74618 585766 0 155 186 156 - 11.5 17.3 50.2 -
case10000_goc 440997 92799 712177 0 - 86 - - - 10.3 - -
case10000_goc__api 440997 92799 712177 0 62 206 - - 6.91 21.1 - -
case10000_goc__sad 440997 92799 712177 0 90 152 - - 7.44 16.4 - -
case8387_pegase 459046 96351 750588 0 139 - 138 - 9.19 - 30.3 -
case8387_pegase__api 459046 96351 750588 0 140 - 140 - 10 - 34.3 -
case8387_pegase__sad 459046 96351 750588 0 134 - 134 - 9.01 - 29.8 -
case9591_goc 495073 102120 811889 0 226 - 229 - 21 - 85.7 -
case9591_goc__api 495073 102120 811889 0 149 159 149 - 15.1 25.5 58.6 -
case9591_goc__sad 495073 102120 811889 0 228 200 228 - 21.3 31.5 86.8 -
case9241_pegase 502110 104741 827419 0 70 - 66 - 6.83 - 20.3 -
case9241_pegase__api 502110 104741 827419 0 67 - 70 - 6.61 - 20.6 -
case9241_pegase__sad 502110 104741 827419 0 70 389 68 - 6.73 58.4 20.5 -
case10192_epigrids 529695 109758 866501 0 96 113 96 - 10.6 17 36 -
case10192_epigrids__api 529695 109758 866501 0 91 113 91 - 10.4 17.4 35.3 -
case10192_epigrids__sad 529695 109758 866501 0 109 - 107 - 11.9 - 41.5 -
case10480_goc 573754 118760 943278 0 164 - 162 - 18.2 - 76.2 -
case10480_goc__api 573754 118760 943278 0 136 217 - - 16.2 40.5 - -
case10480_goc__sad 573754 118760 943278 0 164 242 164 - 18.3 43.6 77.4 -
case13659_pegase 662910 140961 1081042 0 83 - 82 - 11.8 - 42.9 -
case13659_pegase__api 662910 140961 1081042 0 78 - 78 - 11.4 - 39.9 -
case13659_pegase__sad 662910 140961 1081042 0 78 329 - - 11 60 - -
case20758_epigrids 1048059 218151 1709288 0 122 - 123 - 27.5 - 99.1 -
case20758_epigrids__api 1048059 218151 1709288 0 154 - 155 - 39.9 - 150 -
case20758_epigrids__sad 1048059 218151 1709288 0 137 - 135 - 30.3 - 111 -
case19402_goc 1064421 219911 1753874 0 256 - 251 - 45.9 - 221 -
case19402_goc__api 1064421 219911 1753874 0 212 - 216 - 39.3 - 193 -
case19402_goc__sad 1064421 219911 1753874 0 257 - 258 - 46 - 232 -
case30000_goc__api 1195834 249462 1921390 0 150 - - - 24.6 - - -
case24464_goc 1202964 248644 1983519 0 230 - - - 39.4 - - -
case24464_goc__api 1202964 248644 1983519 0 213 - 214 - 33.4 - 149 -
case24464_goc__sad 1202964 248644 1983519 0 - - 224 - - - 169 -
case78484_epigrids 3904758 811998 6389202 0 401 - - - 260 - - -
case78484_epigrids__api 3904758 811998 6389202 0 400 - 401 - 320 - 1.65e+03 -
case78484_epigrids__sad 3904758 811998 6389202 0 408 - - - 263 - - -
Table 6: Solve times and iteration counts for the exponential cone programming problem set
iterations total time (s)
entropy_N_2000 11001 4000 258917 0 23 15 21 10 0.548 0.506 3.05 0.565
entropy_N_4000 22001 8000 1019263 0 24 16 20 9 1.87 2.29 14.2 2.22
entropy_N_6000 33001 12000 2278947 0 26 16 22 9 3.56 6.09 45.1 5.85
entropy_N_8000 44001 16000 4037959 0 27 16 22 10 6.2 12.3 107 12.6
entropy_N_10000 55001 20000 6297376 0 27 16 22 10 9.96 22.6 235 40.1
log_reg_n_2000 94002 34001 1320255 0 57 38 121 37 4.13 3.96 107 5.43
log_reg_n_4000 188002 68001 5205338 0 68 47 - 47 17.4 16.4 - 21.6
log_reg_n_6000 282002 102001 11529263 0 72 49 - 51 40.9 67.7 - 59.8
log_reg_n_8000 376002 136001 20589037 0 73 48 - 50 77.5 135 - 105
log_reg_n_10000 470002 170001 31904473 0 84 54 - 58 150 372 - 336

### v3

@@ Appendix C Detailed benchmark results
case2383wp_k 97600 20393 155950 0 110 - 107 - 1.48 - 2.79 -
case2383wp_k__api 97600 20393 155950 0 21 26 22 33 0.626 0.892 0.868 2.82
case2383wp_k__sad 97600 20393 155950 0 96 - 97 - 1.12 - 2.94 -
case2312_goc 98556 20518 159516 0 47 152 47 - 0.802 4.02 1.58 -
case2312_goc__api 98556 20518 159516 0 65 244 64 - 1.04 6.15 1.71 -
case2312_goc__sad 98556 20518 159516 0 59 191 59 - 0.992 5.08 1.61 -
case2737sop_k 109822 22777 176602 0 74 - 74 - 1.06 - 3.66 -
case2737sop_k__api 109822 22777 176602 0 71 - 71 - 1.01 - 2.37 -
case2737sop_k__sad 109822 22777 176602 0 72 - 71 - 1.03 - 3.07 -
case2736sp_k 110022 22878 176901 0 69 - 74 - 1.24 - 2.15 -
case2736sp_k__api 110022 22878 176901 0 57 - 57 - 0.929 - 1.9 -
case2736sp_k__sad 110022 22878 176901 0 72 - 72 - 1.01 - 4.2 -
case2746wp_k 111106 23320 178556 0 73 - 71 - 1.07 - 2.47 -
case2746wp_k__api 111106 23320 178556 0 33 35 - 34 0.76 1.26 - 2.24
case2746wp_k__sad 111106 23320 178556 0 74 - 75 - 1.07 - 2.26 -
case2746wop_k 111822 23434 179829 0 65 - 65 - 1 - 2.14 -
case2746wop_k__api 111822 23434 179829 0 28 31 25 36 0.913 1.16 1.22 2.44
case2746wop_k__sad 111822 23434 179829 0 64 - 62 - 0.966 - 1.97 -
case3012wp_k 120676 25202 193907 0 86 - 89 - 1.16 - 5.32 -
case3012wp_k__api 120676 25202 193907 0 28 57 - 34 0.77 2.1 - 2.38
case3012wp_k__sad 120676 25202 193907 0 99 - 94 - 1.31 - 2.86 -
case2848_rte 122708 25858 197228 0 81 - 81 - 1.3 - 3.4 -
case2848_rte__api 122708 25858 197228 0 76 218 75 - 1.35 6.1 3.7 -
case2848_rte__sad 122708 25858 197228 0 70 217 70 - 1.25 6 3.23 -
case2868_rte 123912 26164 198869 0 84 228 86 - 1.41 6.72 8.48 -
case2868_rte__api 123912 26164 198869 0 72 258 77 - 1.55 7.62 3.84 -
case2868_rte__sad 123912 26164 198869 0 79 287 79 - 1.32 8.18 6.56 -
case3120sp_k 124354 25856 199827 0 82 - 95 - 1.21 - 4.85 -
case3120sp_k__api 124354 25856 199827 0 96 - 93 - 1.32 - 5.08 -
case2853_sdet 128886 27445 206896 0 103 - 105 - 1.29 - 6.31 -
case2853_sdet__api 128886 27445 206896 0 120 - 116 - 1.41 - 3.45 -
case2853_sdet__sad 128886 27445 206896 0 102 278 102 - 1.27 8.68 3.01 -
case3022_goc 134780 28060 217434 0 41 176 - 140 0.984 5.86 - 9.95
case3022_goc__api 134780 28060 217434 0 55 203 54 170 1.1 6.8 2.2 12.3
case3022_goc__sad 134780 28060 217434 0 41 177 41 141 0.971 5.76 1.74 9.79
case3375wp_k 139126 29112 223999 0 86 - 85 - 1.26 - 3.05 -
case3375wp_k__api 139126 29112 223999 0 161 - 147 - 1.86 - 4.82 -
case3375wp_k__sad 139126 29112 223999 0 104 - 93 - 1.4 - 3.48 -
case2869_pegase 143608 30153 236217 0 53 236 53 - 1.25 8.46 4.9 -
case2869_pegase__api 143608 30153 236217 0 50 262 50 - 1.24 9.31 6.44 -
case2869_pegase__sad 143608 30153 236217 0 48 278 48 - 1.22 9.66 2.99 -
case2742_goc 144110 29856 236467 0 109 118 113 - 2.31 4.81 14.4 -
case2742_goc__api 144110 29856 236467 0 101 128 103 - 2.15 5.06 7.2 -
case2742_goc__sad 144110 29856 236467 0 106 118 108 - 2.32 4.67 7.47 -
case4661_sdet 198498 41599 320728 0 139 - 140 - 2.7 - 16.8 -
case4661_sdet__api 198498 41599 320728 0 126 - 124 - 2.59 - 10.7 -
case4661_sdet__sad 198498 41599 320728 0 140 322 138 - 2.71 18.6 8.3 -
case3970_goc 205819 42789 337328 0 140 156 139 - 3.79 9.51 12 -
case3970_goc__api 205819 42789 337328 0 114 137 116 - 3.59 8.41 12.3 -
case3970_goc__sad 205819 42789 337328 0 147 168 141 - 3.23 10.2 12 -
case4020_goc 216455 44877 355706 0 187 - 180 - 4.64 - 18.3 -
case4020_goc__api 216455 44877 355706 0 98 142 101 - 3.13 8.86 21.7 -
case4020_goc__sad 216455 44877 355706 0 180 - 187 - 4.57 - 20.6 -
case4917_goc 218213 45522 352833 0 52 166 52 147 1.7 8.82 4.82 16.5
case4917_goc__api 218213 45522 352833 0 62 256 62 - 1.86 13.4 5.98 -
case4917_goc__sad 218213 45522 352833 0 48 183 - 169 1.69 9.8 - 18.7
case4601_goc 225588 46885 368445 0 180 192 180 - 3.84 11.7 19.8 -
case4601_goc__api 225588 46885 368445 0 109 148 109 156 3.03 8.96 14.4 19.3
case4601_goc__sad 225588 46885 368445 0 179 191 178 - 4.56 12.2 15 -
case4837_goc 240160 49855 392884 0 134 154 138 - 3.82 10.3 36.1 -
case4837_goc__api 240160 49855 392884 0 123 177 123 - 3.73 12.1 13.7 -
case4837_goc__sad 240160 49855 392884 0 133 162 131 - 3.88 11 17.5 -
case4619_goc 254753 52616 419210 0 147 151 145 160 4.51 10.7 19.9 22.2
case4619_goc__api 254753 52616 419210 0 129 161 131 - 4.01 11.7 15.8 -
case4619_goc__sad 254753 52616 419210 0 148 151 150 - 5.15 11 19.2 -
case5658_epigrids 282180 58620 461724 0 146 177 143 - 5.53 12.9 19.7 -
case5658_epigrids__api 282180 58620 461724 0 163 203 162 - 5.45 14.6 25.3 -
case5658_epigrids__sad 282180 58620 461724 0 163 217 157 - 4.8 15.7 20.4 -
case6468_rte 286248 59396 463599 0 103 250 98 - 3.01 15.8 12.4 -
case6468_rte__api 286248 59396 463599 0 88 - 90 - 3.59 - 15.2 -
case6468_rte__sad 286248 59396 463599 0 98 255 99 - 3.08 16.4 19.7 -
case6470_rte 287806 60144 465757 0 113 321 111 - 3.29 20.5 26.4 -
case6470_rte__api 287806 60144 465757 0 85 327 85 - 2.95 21.1 15.5 -
case6470_rte__sad 287806 60144 465757 0 104 354 100 - 3.66 22.2 14.4 -
case6495_rte 288050 60099 465939 0 114 323 114 - 3.37 20.4 16.4 -
case6495_rte__api 288050 60099 465939 0 - 343 98 - - 21.4 12.4 -
case6495_rte__sad 288050 60099 465939 0 112 344 108 - 3.25 22.1 25.2 -
case6515_rte 288710 60239 466922 0 97 294 95 - 3.19 18.3 12.6 -
case6515_rte__api 288710 60239 466922 0 85 341 81 - 2.97 21.4 21.4 -
case6515_rte__sad 288710 60239 466922 0 84 324 88 - 2.79 20.4 12.1 -
case7336_epigrids 358450 74618 585766 0 155 195 153 - 6.51 17.2 35.9 -
case7336_epigrids__api 358450 74618 585766 0 152 206 151 - 6.1 18.4 32.8 -
case7336_epigrids__sad 358450 74618 585766 0 156 188 156 - 6.56 17.8 58.6 -
case10000_goc 440997 92799 712177 0 71 87 - - 3.58 9.49 - -
case10000_goc__api 440997 92799 712177 0 62 205 63 - 3.97 20.4 15.6 -
case10000_goc__sad 440997 92799 712177 0 89 152 - - 4.13 16 - -
case8387_pegase 459046 96351 750588 0 137 - 138 - 4.8 - 22.3 -
case8387_pegase__api 459046 96351 750588 0 141 - 141 - 5.08 - 25.9 -
case8387_pegase__sad 459046 96351 750588 0 - - 142 - - - 34.9 -
case9591_goc 495073 102120 811889 0 229 206 226 - 12 32.7 69.9 -
case9591_goc__api 495073 102120 811889 0 149 162 149 190 7.92 26.3 41.6 49.3
case9591_goc__sad 495073 102120 811889 0 230 - 227 - 10.8 - 78.6 -
case9241_pegase 502110 104741 827419 0 72 - 67 - 5.29 - 14.9 -
case9241_pegase__api 502110 104741 827419 0 69 - - - 4.02 - - -
case9241_pegase__sad 502110 104741 827419 0 72 - 68 - 4.26 - 21.7 -
case10192_epigrids 529695 109758 866501 0 94 113 95 - 5.74 16.5 29.5 -
case10192_epigrids__api 529695 109758 866501 0 93 115 - - 5.92 16.7 - -
case10192_epigrids__sad 529695 109758 866501 0 110 - 114 - 6.36 - 33.2 -
case10480_goc 573754 118760 943278 0 160 235 161 - 9.62 42.4 53.5 -
case10480_goc__api 573754 118760 943278 0 139 213 - - 8.91 39.3 - -
case10480_goc__sad 573754 118760 943278 0 162 - 164 - 11 - 54.3 -
case13659_pegase 662910 140961 1081042 0 80 - 82 - 6.53 - 30.8 -
case13659_pegase__api 662910 140961 1081042 0 78 - 78 - 6.35 - 29.2 -
case13659_pegase__sad 662910 140961 1081042 0 78 331 77 - 6.29 59.7 30.5 -
case20758_epigrids 1048059 218151 1709288 0 123 - - - 12.4 - - -
case20758_epigrids__api 1048059 218151 1709288 0 153 - 150 - 16 - 112 -
case20758_epigrids__sad 1048059 218151 1709288 0 136 - 136 - 15.6 - 79.9 -
case19402_goc 1064421 219911 1753874 0 259 - 254 - 21.8 - 168 -
case19402_goc__api 1064421 219911 1753874 0 215 - 215 - 19.3 - 165 -
case19402_goc__sad 1064421 219911 1753874 0 253 - 253 - 21.5 - 175 -
case30000_goc__api 1195834 249462 1921390 0 146 - - - 13.1 - - -
case24464_goc 1202964 248644 1983519 0 233 - 219 - 21.2 - 133 -
case24464_goc__api 1202964 248644 1983519 0 227 - 210 - 16.9 - 108 -
case24464_goc__sad 1202964 248644 1983519 0 226 - 221 - 17.9 - 124 -
case78484_epigrids 3904758 811998 6389202 0 413 - 410 - 118 - 1.02e+03 -
case78484_epigrids__api 3904758 811998 6389202 0 396 - - - 138 - - -
case78484_epigrids__sad 3904758 811998 6389202 0 405 - 405 - 103 - 981 -